OpenStudy (anonymous):

what is the slope of the line passing through the points (2 7) and (–1 3) ? A. 2/7 B. 3/4 C. 4/3 D. 1/3

OpenStudy (shamil98):

y -y1 = m(x-x1)

OpenStudy (jdoe0001):

\(\large slope = m= \cfrac{rise}{run} \implies \cfrac{y_2-y_1}{x_2-x_1}\)

OpenStudy (shamil98):

But, first find the slope.

OpenStudy (shamil98):

y2-y1/x2-x1 = m

OpenStudy (anonymous):

so 7-3 = m(2--1) ?